\begin{thebibliography}{46}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Bahdanau et~al.(2015)Bahdanau, Cho, and Bengio}]{bahdanau2014neural}
Dzmitry Bahdanau, Kyunghyun Cho, and Yoshua Bengio. 2015.
\newblock \href {http://arxiv.org/abs/1409.0473} {Neural machine translation by
  jointly learning to align and translate}.
\newblock In \emph{3rd International Conference on Learning Representations,
  {ICLR} 2015, San Diego, CA, USA, May 7-9, 2015, Conference Track
  Proceedings}.

\bibitem[{Caswell et~al.(2019)Caswell, Chelba, and
  Grangier}]{caswell2019tagged}
Isaac Caswell, Ciprian Chelba, and David Grangier. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-5206} {Tagged
  back-translation}.
\newblock In \emph{Proceedings of the Fourth Conference on Machine Translation
  (Volume 1: Research Papers)}, pages 53--63, Florence, Italy. Association for
  Computational Linguistics.

\bibitem[{Chen et~al.(2020)Chen, Gan, Cheng, Liu, and Liu}]{chen2020distilling}
Yen-Chun Chen, Zhe Gan, Yu~Cheng, Jingzhou Liu, and Jingjing Liu. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.705} {Distilling
  knowledge learned in {BERT} for text generation}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 7893--7905, Online. Association for
  Computational Linguistics.

\bibitem[{Conneau et~al.(2020)Conneau, Khandelwal, Goyal, Chaudhary, Wenzek,
  Guzm{\'a}n, Grave, Ott, Zettlemoyer, and Stoyanov}]{conneau2019unsupervised}
Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume
  Wenzek, Francisco Guzm{\'a}n, Edouard Grave, Myle Ott, Luke Zettlemoyer, and
  Veselin Stoyanov. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.747} {Unsupervised
  cross-lingual representation learning at scale}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 8440--8451, Online. Association for
  Computational Linguistics.

\bibitem[{Conneau and Lample(2019)}]{lample2019cross}
Alexis Conneau and Guillaume Lample. 2019.
\newblock \href
  {https://proceedings.neurips.cc/paper/2019/hash/c04c19c2c2474dbf5f7ac4372c5b9af1-Abstract.html}
  {Cross-lingual language model pretraining}.
\newblock In \emph{Advances in Neural Information Processing Systems 32: Annual
  Conference on Neural Information Processing Systems 2019, NeurIPS 2019,
  December 8-14, 2019, Vancouver, BC, Canada}, pages 7057--7067.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and Toutanova}]{Devlin:2018uk}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Ding et~al.(2021{\natexlab{a}})Ding, Wang, Liu, Wong, Tao, and
  Tu}]{ding-etal-2021-progressive}
Liang Ding, Longyue Wang, Xuebo Liu, Derek~F. Wong, Dacheng Tao, and Zhaopeng
  Tu. 2021{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2021.findings-acl.247}
  {Progressive multi-granularity training for non-autoregressive translation}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021}, pages 2797--2803, Online. Association for Computational
  Linguistics.

\bibitem[{Ding et~al.(2021{\natexlab{b}})Ding, Wang, Liu, Wong, Tao, and
  Tu}]{ding-etal-2021-rejuvenating}
Liang Ding, Longyue Wang, Xuebo Liu, Derek~F. Wong, Dacheng Tao, and Zhaopeng
  Tu. 2021{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.266} {Rejuvenating
  low-frequency words: Making the most of parallel data in non-autoregressive
  translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 3431--3441,
  Online. Association for Computational Linguistics.

\bibitem[{Ding et~al.(2021{\natexlab{c}})Ding, Wu, and
  Tao}]{ding-etal-2021-improving}
Liang Ding, Di~Wu, and Dacheng Tao. 2021{\natexlab{c}}.
\newblock \href {https://arxiv.org/abs/2109.07780} {Improving neural machine
  translation by bidirectional training}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}. Association for Computational
  Linguistics.

\bibitem[{Edunov et~al.(2018)Edunov, Ott, Auli, and
  Grangier}]{edunov-etal-2018-understanding}
Sergey Edunov, Myle Ott, Michael Auli, and David Grangier. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1045} {Understanding
  back-translation at scale}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 489--500, Brussels, Belgium. Association
  for Computational Linguistics.

\bibitem[{Edunov et~al.(2020)Edunov, Ott, Ranzato, and
  Auli}]{edunov2020evaluation}
Sergey Edunov, Myle Ott, Marc{'}Aurelio Ranzato, and Michael Auli. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.253} {On the
  evaluation of machine translation systems trained with back-translation}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 2836--2846, Online. Association for
  Computational Linguistics.

\bibitem[{Fadaee and Monz(2018)}]{fadaee2018back}
Marzieh Fadaee and Christof Monz. 2018.
\newblock \href {https://doi.org/10.18653/v1/D18-1040} {Back-translation
  sampling by targeting difficult words in neural machine translation}.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 436--446, Brussels, Belgium. Association
  for Computational Linguistics.

\bibitem[{Gehring et~al.(2017)Gehring, Auli, Grangier, Yarats, and
  Dauphin}]{DBLP:journals/corr/GehringAGYD17}
Jonas Gehring, Michael Auli, David Grangier, Denis Yarats, and Yann~N. Dauphin.
  2017.
\newblock \href {http://proceedings.mlr.press/v70/gehring17a.html}
  {Convolutional sequence to sequence learning}.
\newblock In \emph{Proceedings of the 34th International Conference on Machine
  Learning, {ICML} 2017, Sydney, NSW, Australia, 6-11 August 2017}, volume~70
  of \emph{Proceedings of Machine Learning Research}, pages 1243--1252. {PMLR}.

\bibitem[{Graham et~al.(2020)Graham, Haddow, and
  Koehn}]{graham-etal-2020-statistical}
Yvette Graham, Barry Haddow, and Philipp Koehn. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.6} {Statistical
  power and translationese in machine translation evaluation}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 72--81, Online. Association for
  Computational Linguistics.

\bibitem[{He et~al.(2020)He, Gu, Shen, and Ranzato}]{He2020Revisiting}
Junxian He, Jiatao Gu, Jiajun Shen, and Marc'Aurelio Ranzato. 2020.
\newblock \href {https://openreview.net/forum?id=SJgdnAVKDH} {Revisiting
  self-training for neural sequence generation}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Huang et~al.(2021)Huang, Liu, Wang, Wang, Li, Tu, Huang, and
  Shi}]{transmart}
Guoping Huang, Lemao Liu, Xing Wang, Longyue Wang, Huayang Li, Zhaopeng Tu,
  Chengyan Huang, and Shuming Shi. 2021.
\newblock \href {http://arxiv.org/abs/2105.13072} {Transmart: {A} practical
  interactive machine translation system}.
\newblock \emph{CoRR}, abs/2105.13072.

\bibitem[{Jiao et~al.(2021)Jiao, Wang, Tu, Shi, Lyu, and
  King}]{jiao-etal-2021-self}
Wenxiang Jiao, Xing Wang, Zhaopeng Tu, Shuming Shi, Michael Lyu, and Irwin
  King. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.221} {Self-training
  sampling with monolingual data uncertainty for neural machine translation}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 2840--2850,
  Online. Association for Computational Linguistics.

\bibitem[{Leong et~al.(2021)Leong, Liu, Wong, and Chao}]{Leong-etal-exploiting}
Chongman Leong, Xuebo Liu, Derek~F. Wong, and Lidia~S. Chao. 2021.
\newblock \href {https://doi.org/10.1109/TASLP.2021.3105798} {Exploiting
  translation model for parallel corpus mining}.
\newblock \emph{IEEE/ACM Transactions on Audio, Speech, and Language
  Processing}, 29:2829--2839.

\bibitem[{Lewis et~al.(2020)Lewis, Liu, Goyal, Ghazvininejad, Mohamed, Levy,
  Stoyanov, and Zettlemoyer}]{bart2020}
Mike Lewis, Yinhan Liu, Naman Goyal, Marjan Ghazvininejad, Abdelrahman Mohamed,
  Omer Levy, Veselin Stoyanov, and Luke Zettlemoyer. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.703} {{BART}:
  Denoising sequence-to-sequence pre-training for natural language generation,
  translation, and comprehension}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 7871--7880, Online. Association for
  Computational Linguistics.

\bibitem[{Lin et~al.(2020)Lin, Pan, Wang, Qiu, Feng, Zhou, and Li}]{lin2020pre}
Zehui Lin, Xiao Pan, Mingxuan Wang, Xipeng Qiu, Jiangtao Feng, Hao Zhou, and
  Lei Li. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.210} {Pre-training
  multilingual neural machine translation by leveraging alignment information}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 2649--2663, Online. Association
  for Computational Linguistics.

\bibitem[{Liu et~al.(2020{\natexlab{a}})Liu, Lai, Wong, and
  Chao}]{liu-etal-2020-norm}
Xuebo Liu, Houtim Lai, Derek~F. Wong, and Lidia~S. Chao. 2020{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.41} {Norm-based
  curriculum learning for neural machine translation}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 427--436, Online. Association for
  Computational Linguistics.

\bibitem[{Liu et~al.(2021{\natexlab{a}})Liu, Wang, Wong, Ding, Chao, Shi, and
  Tu}]{liu-etal-2021-copying}
Xuebo Liu, Longyue Wang, Derek~F. Wong, Liang Ding, Lidia~S. Chao, Shuming Shi,
  and Zhaopeng Tu. 2021{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2021.findings-acl.373} {On the
  copying behaviors of pre-training for neural machine translation}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021}, pages 4265--4275, Online. Association for Computational
  Linguistics.

\bibitem[{Liu et~al.(2021{\natexlab{b}})Liu, Wang, Wong, Ding, Chao, and
  Tu}]{liu2021understanding}
Xuebo Liu, Longyue Wang, Derek~F. Wong, Liang Ding, Lidia~S. Chao, and Zhaopeng
  Tu. 2021{\natexlab{b}}.
\newblock \href {https://openreview.net/forum?id=n1HD8M6WGn} {Understanding and
  improving encoder layer fusion in sequence-to-sequence learning}.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{Liu et~al.(2020{\natexlab{b}})Liu, Gu, Goyal, Li, Edunov,
  Ghazvininejad, Lewis, and Zettlemoyer}]{Liu:2020mbart}
Yinhan Liu, Jiatao Gu, Naman Goyal, Xian Li, Sergey Edunov, Marjan
  Ghazvininejad, Mike Lewis, and Luke Zettlemoyer. 2020{\natexlab{b}}.
\newblock \href {https://doi.org/10.1162/tacl_a_00343} {Multilingual denoising
  pre-training for neural machine translation}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  8:726--742.

\bibitem[{Marie et~al.(2020)Marie, Rubino, and Fujita}]{marie2020tagged}
Benjamin Marie, Raphael Rubino, and Atsushi Fujita. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.532} {Tagged
  back-translation revisited: Why does it really work?}
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 5990--5997, Online. Association for
  Computational Linguistics.

\bibitem[{Neubig et~al.(2019)Neubig, Dou, Hu, Michel, Pruthi, and
  Wang}]{neubig-etal-2019-compare}
Graham Neubig, Zi-Yi Dou, Junjie Hu, Paul Michel, Danish Pruthi, and Xinyi
  Wang. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-4007} {compare-mt: A tool for
  holistic comparison of language generation systems}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics (Demonstrations)},
  pages 35--41, Minneapolis, Minnesota. Association for Computational
  Linguistics.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni2002bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock \href {https://doi.org/10.3115/1073083.1073135} {{B}leu: a method for
  automatic evaluation of machine translation}.
\newblock In \emph{Proceedings of the 40th Annual Meeting of the Association
  for Computational Linguistics}, pages 311--318, Philadelphia, Pennsylvania,
  USA. Association for Computational Linguistics.

\bibitem[{Post(2018)}]{post-2018-call}
Matt Post. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-6319} {A call for clarity in
  reporting {BLEU} scores}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 186--191, Brussels, Belgium. Association for
  Computational Linguistics.

\bibitem[{Rothe et~al.(2020)Rothe, Narayan, and Severyn}]{rothe2020leveraging}
Sascha Rothe, Shashi Narayan, and Aliaksei Severyn. 2020.
\newblock \href {https://doi.org/10.1162/tacl_a_00313} {Leveraging pre-trained
  checkpoints for sequence generation tasks}.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  8:264--280.

\bibitem[{Sennrich et~al.(2016{\natexlab{a}})Sennrich, Haddow, and
  Birch}]{DBLP:journals/corr/SennrichHB16}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/W16-2323} {{E}dinburgh neural
  machine translation systems for {WMT} 16}.
\newblock In \emph{Proceedings of the First Conference on Machine Translation:
  Volume 2, Shared Task Papers}, pages 371--376, Berlin, Germany. Association
  for Computational Linguistics.

\bibitem[{Sennrich et~al.(2016{\natexlab{b}})Sennrich, Haddow, and
  Birch}]{sennrich2016improving}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/P16-1009} {Improving neural
  machine translation models with monolingual data}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 86--96, Berlin,
  Germany. Association for Computational Linguistics.

\bibitem[{Sennrich et~al.(2016{\natexlab{c}})Sennrich, Haddow, and
  Birch}]{DBLP:journals/corr/SennrichHB15}
Rico Sennrich, Barry Haddow, and Alexandra Birch. 2016{\natexlab{c}}.
\newblock \href {https://doi.org/10.18653/v1/P16-1162} {Neural machine
  translation of rare words with subword units}.
\newblock In \emph{Proceedings of the 54th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 1715--1725,
  Berlin, Germany. Association for Computational Linguistics.

\bibitem[{Snover et~al.(2006)Snover, Dorr, Schwartz, Micciulla, and
  Makhoul}]{snover2006study}
Matthew Snover, Bonnie Dorr, Richard Schwartz, Linnea Micciulla, and John
  Makhoul. 2006.
\newblock \href {http://www.cs.umd.edu/~snover/pub/amta06/ter_amta.pdf} {A
  study of translation edit rate with targeted human annotation}.
\newblock In \emph{Proceedings of association for machine translation in the
  Americas}, volume 200. Citeseer.

\bibitem[{Song et~al.(2019)Song, Tan, Qin, Lu, and Liu}]{song2019mass}
Kaitao Song, Xu~Tan, Tao Qin, Jianfeng Lu, and Tie{-}Yan Liu. 2019.
\newblock \href {http://proceedings.mlr.press/v97/song19d.html} {{MASS:} masked
  sequence to sequence pre-training for language generation}.
\newblock In \emph{Proceedings of the 36th International Conference on Machine
  Learning, {ICML} 2019, 9-15 June 2019, Long Beach, California, {USA}},
  volume~97 of \emph{Proceedings of Machine Learning Research}, pages
  5926--5936. {PMLR}.

\bibitem[{Tang et~al.(2019)Tang, Sennrich, and Nivre}]{tang2019encoders}
Gongbo Tang, Rico Sennrich, and Joakim Nivre. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1149} {Encoders help you
  disambiguate word senses in neural machine translation}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 1429--1435, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Tang et~al.(2020)Tang, Tran, Li, Chen, Goyal, Chaudhary, Gu, and
  Fan}]{tang2020multilingual}
Yuqing Tang, Chau Tran, Xian Li, Peng-Jen Chen, Naman Goyal, Vishrav Chaudhary,
  Jiatao Gu, and Angela Fan. 2020.
\newblock \href {https://arxiv.org/abs/2008.00401} {Multilingual translation
  with extensible multilingual pretraining and finetuning}.
\newblock \emph{arXiv preprint arXiv:2008.00401}.

\bibitem[{Tran et~al.(2020)Tran, Tang, Li, and Gu}]{tran2020cross}
Chau Tran, Yuqing Tang, Xian Li, and Jiatao Gu. 2020.
\newblock \href
  {https://proceedings.neurips.cc/paper/2020/file/1763ea5a7e72dd7ee64073c2dda7a7a8-Paper.pdf}
  {Cross-lingual retrieval for iterative self-supervised training}.
\newblock In \emph{Advances in Neural Information Processing Systems},
  volume~33, pages 2207--2219. Curran Associates, Inc.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{DBLP:journals/corr/VaswaniSPUJGKP17}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N. Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017.
\newblock \href
  {https://proceedings.neurips.cc/paper/2017/hash/3f5ee243547dee91fbd053c1c4a845aa-Abstract.html}
  {Attention is all you need}.
\newblock In \emph{Advances in Neural Information Processing Systems 30: Annual
  Conference on Neural Information Processing Systems 2017, December 4-9, 2017,
  Long Beach, CA, {USA}}, pages 5998--6008.

\bibitem[{Wang et~al.(2019)Wang, Liu, Wang, Luan, and Sun}]{wang2019improving}
Shuo Wang, Yang Liu, Chao Wang, Huanbo Luan, and Maosong Sun. 2019.
\newblock \href {https://doi.org/10.18653/v1/D19-1073} {Improving
  back-translation with uncertainty-based confidence estimation}.
\newblock In \emph{Proceedings of the 2019 Conference on Empirical Methods in
  Natural Language Processing and the 9th International Joint Conference on
  Natural Language Processing (EMNLP-IJCNLP)}, pages 791--802, Hong Kong,
  China. Association for Computational Linguistics.

\bibitem[{Wang et~al.(2021)Wang, Tu, Tan, Shi, Sun, and
  Liu}]{wang-etal-2021-language-coverage}
Shuo Wang, Zhaopeng Tu, Zhixing Tan, Shuming Shi, Maosong Sun, and Yang Liu.
  2021.
\newblock \href {https://doi.org/10.18653/v1/2021.findings-acl.422} {On the
  language coverage bias for neural machine translation}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021}, pages 4778--4790, Online. Association for Computational
  Linguistics.

\bibitem[{Yang et~al.(2020{\natexlab{a}})Yang, Wang, Zhou, Zhao, Zhang, Yu, and
  Li}]{yang2020towards}
Jiacheng Yang, Mingxuan Wang, Hao Zhou, Chengqi Zhao, Weinan Zhang, Yong Yu,
  and Lei Li. 2020{\natexlab{a}}.
\newblock \href {https://ojs.aaai.org//index.php/AAAI/article/view/6479}
  {Towards making the most of bert in neural machine translation}.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, volume~34, pages 9378--9385.

\bibitem[{Yang et~al.(2020{\natexlab{b}})Yang, Hu, Han, Huang, and
  Ju}]{yang2020csp}
Zhen Yang, Bojie Hu, Ambyera Han, Shen Huang, and Qi~Ju. 2020{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.208}
  {{CSP}:code-switching pre-training for neural machine translation}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 2624--2636, Online. Association
  for Computational Linguistics.

\bibitem[{Zhan et~al.(2021)Zhan, Liu, Wong, and Chao}]{Zhan_Liu_Wong_Chao_2021}
Runzhe Zhan, Xuebo Liu, Derek~F. Wong, and Lidia~S. Chao. 2021.
\newblock \href {https://ojs.aaai.org/index.php/AAAI/article/view/17683}
  {Meta-curriculum learning for domain adaptation in neural machine
  translation}.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, pages 14310--14318.

\bibitem[{Zhang and Zong(2016)}]{zhang-zong-2016-exploiting}
Jiajun Zhang and Chengqing Zong. 2016.
\newblock \href {https://doi.org/10.18653/v1/D16-1160} {Exploiting source-side
  monolingual data in neural machine translation}.
\newblock In \emph{Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 1535--1545, Austin, Texas. Association
  for Computational Linguistics.

\bibitem[{Zhang and Toral(2019)}]{zhang2019effect}
Mike Zhang and Antonio Toral. 2019.
\newblock \href {https://doi.org/10.18653/v1/W19-5208} {The effect of
  translationese in machine translation test sets}.
\newblock In \emph{Proceedings of the Fourth Conference on Machine Translation
  (Volume 1: Research Papers)}, pages 73--81, Florence, Italy. Association for
  Computational Linguistics.

\bibitem[{Zhu et~al.(2020)Zhu, Xia, Wu, He, Qin, Zhou, Li, and
  Liu}]{Zhu2020Incorporating}
Jinhua Zhu, Yingce Xia, Lijun Wu, Di~He, Tao Qin, Wengang Zhou, Houqiang Li,
  and Tie{-}Yan Liu. 2020.
\newblock \href {https://openreview.net/forum?id=Hyl7ygStwB} {Incorporating
  {BERT} into neural machine translation}.
\newblock In \emph{8th International Conference on Learning Representations,
  {ICLR} 2020, Addis Ababa, Ethiopia, April 26-30, 2020}. OpenReview.net.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{emnlp2021}
\usepackage{times}
\usepackage{latexsym}
\usepackage{color}
\usepackage{multirow}
\usepackage{forest}
\usepackage{tikz}
\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage{amsmath}
\usepackage{graphics}
\usepackage{subfigure}
\usepackage{mwe}
\usepackage{amssymb}
\usepackage{bm}
\usepackage{amsmath}
\usepackage{makecell}
\usepackage{tabularx}
\usepackage{url}
\usepackage{pgfplots}
\usepackage{CJKutf8}
\pgfplotsset{width=10cm,compat=1.9}
\usepackage{booktabs}
\usepackage{dsfont}
\usepackage{arydshln}
\usepackage{enumitem}

\usepackage{microtype}

\definecolor{electricviolet}{rgb}{0.56, 0.0, 1.0}

\usetikzlibrary{arrows.meta,calc,decorations.markings,math,arrows.meta,patterns}
\usetikzlibrary{decorations.pathreplacing}
\usetikzlibrary{matrix}

\title{On the Complementarity between Pre-Training and Back-Translation \\ for Neural Machine Translation}

\author{Xuebo Liu$^{1}$\thanks{~~Work was done when Xuebo Liu and Liang Ding were interning at Tencent AI Lab.} , Longyue Wang$^{2}$, Derek F. Wong$^{1}$, Liang Ding$^{3}$, \\ \bf Lidia S. Chao$^{1}$, Shuming Shi$^{2}$ and Zhaopeng Tu$^{2}$\\
  $^{1}$NLP$^2$CT Lab, Department of Computer and Information Science, 
  University of Macau \\
  $^{2}$Tencent AI Lab~~~~$^{3}$The University of Sydney \\
  $^{1}${\tt nlp2ct.xuebo@gmail.com, \{derekfw,lidiasc\}@um.edu.com} \\
  $^{2}${\tt \{vinnylywang,shumingshi,zptu\}@tencent.com} \\
  $^{3}${\tt ldin3097@sydney.edu.au}}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Pre-training (PT) and back-translation (BT) are two simple and powerful methods to utilize monolingual data for improving the model performance of neural machine translation (NMT).
This paper takes the first step to investigate the complementarity between PT and BT.
We introduce two probing tasks for PT and BT respectively and find that PT mainly contributes to the encoder module while BT brings more benefits to the decoder.
Experimental results show that PT and BT are nicely complementary to each other, establishing state-of-the-art performances on the WMT16 English-Romanian and English-Russian benchmarks.
Through extensive analyses on sentence originality and word frequency, we also demonstrate that combining Tagged BT with PT is more helpful to their complementarity, leading to better translation quality.
Source code is freely available at \url{https://github.com/SunbowLiu/PTvsBT}.
\end{abstract}

\section{Introduction}
Neural machine translation (\citealp[NMT;][]{bahdanau2014neural,DBLP:journals/corr/GehringAGYD17,DBLP:journals/corr/VaswaniSPUJGKP17}) models are data-hungry and their performances are highly dependent upon the quantity and quality of labeled data, which are expensive and scarce resources~\citep{Leong-etal-exploiting}.
This motivates the research line of exploiting unlabeled monolingual data for boosting the model performance of NMT.
Due to simplicity and effectiveness, pre-training (\citealp[PT;][]{Devlin:2018uk,song2019mass}) and back-translation (\citealp[BT;][]{sennrich2016improving}) are two widely-used techniques for NMT, by leveraging a large amount of monolingual data.

While empirically successful, the understandings of PT and BT are still limited at best. 
Several attempts have been made to better understand them at the data level, e.g. exploring different kinds of noises for the source data~\citep{edunov-etal-2018-understanding,bart2020}. 
However, there are few understandings at the model level that how PT and BT affect the internal module (e.g. encoder and decoder) of NMT models. 
As recent studies start to combine PT and BT for better model performance~\citep{lample2019cross,Liu:2020mbart,ding-etal-2021-improving}, there is a pressing need to broaden the understandings of them. 

To this end, we introduce two probing tasks to investigate the effects of PT and BT on the encoder and decoder modules, respectively. 
We find that PT mainly contributes to the encoder module while BT brings more benefits to the decoder module. 
This provides a good explanation for the performance improvement of simply combining PT and BT. 
Motivated by this finding, we explore a better combination method by leveraging {\it Tagged BT}~\citep{caswell2019tagged}. 
Experiments conducted on the WMT16 English-Romanian and English-Russian benchmarks show that PT can nicely co-work with BT, leading to state-of-the-art model performances. 
Extensive analyses show that the tagging mechanism is helpful for enhancing the complementarity between PT and BT by improving the translation of source-original sentences and low-frequency words. 

Our {\bf main contributions} are as follows:
\begin{itemize}
\item We design two probing tasks to investigate the impact of PT and BT on NMT models.

\item We empirically demonstrate the complementarity between PT and BT.

\item We show that Tagged BT further improves the complementarity between PT and BT.
\end{itemize}

\section{Preliminaries}

\subsection{Background}

\paragraph{Pre-Training for NMT}
Self-supervised PT~\citep{Devlin:2018uk,song2019mass}, which can acquire knowledge from unlabeled monolingual data, has shown its effectiveness in improving the model performance of NMT, especially for those language pairs with smaller parallel corpora~\citep{lample2019cross}.

The first research line treats pre-trained models as external knowledge to guidance NMT to learn better representations~\citep{yang2020towards,Zhu2020Incorporating} and predictions~\citep{chen2020distilling}.
These methods are effective but costly since the NMT architecture needed to be elaborately designed.
Another research line is directly taking the weights of pre-trained models to initialize NMT models, which is easy to use and advancing the state-of-the-art~\citep{rothe2020leveraging,bart2020}.
In this paper, we treat pre-trained mBART~\citep{Liu:2020mbart} as our testbed for parameter initialization, whose benefits have been sufficiently validated~\citep{tran2020cross,tang2020multilingual,liu-etal-2021-copying} by multiple translation directions.

In general, previous studies focus on designing novel architectures~\citep{song2019mass} and artificial noises for source sentences~\citep{lin2020pre,yang2020csp} but are still unclear why pre-training can boost the model performance of NMT, which is this paper aims to investigate.

\paragraph{Back-Translation for NMT}
BT is an alternative to leverage monolingual data for NMT~\citep{sennrich2016improving}.
It first trains a reversed NMT model for translating target-side monolingual data into synthetic parallel data, and then complements them with the original parallel data to train the desired NMT model.
To improve BT, previous works put attention to the importance of diversity and complexity in synthetic data, showing that adding symbols (e.g., noises and tags) to the back-translated source can help NMT distinguish the data from various sources and learn better representations~\citep{fadaee2018back,wang2019improving,edunov-etal-2018-understanding,caswell2019tagged,marie2020tagged}. 
The claims and understandings from these works are chiefly at the data-level rather than the model-level.

There also exists some works that combine PT and BT to further boost the model performance~\citep{conneau2019unsupervised,song2019mass,Liu:2020mbart}.
However, the relation between BT and PT has not been fully studied.
In this paper, we take the first step to understand BT and PT at the model-level and improve the complementarity between PT and BT.

\subsection{Experimental Setup}
\label{sec:setup}
\paragraph{Data}
We conducted experiments on the WMT16 English-Romanian (En-Ro) and English-Russia (En-Ru) translation tasks, which are widely-used benchmarks of data augmentation methods for NMT.
The training/validation/test sets of the En-Ro include 612K/2K/2K sentence pairs, while those of En-Ru include 2M/3K/3K pairs.
Towards better reproducibility, we directly used the BT data provided by~\citet{DBLP:journals/corr/SennrichHB16}\footnote{\url{http://data.statmt.org/rsennrich/wmt16_backtranslations}}, consisting of 2.3M synthetic data for the En-Ro and 2.0M data for the En-Ru.
All the data are tokenized and split into sub-words~\citep{DBLP:journals/corr/SennrichHB15} by the mBART tokenizer~\citep{Liu:2020mbart}.

\paragraph{Setting}
To make a fair comparison, all the model architectures and parameters are the same as the pre-trained {\tt mBART.cc25}.\footnote{\url{https://github.com/pytorch/fairseq/tree/master/examples/mbart}}
The NMT model augmented with PT directly uses the mBART weights for parameter initialization, while the other models randomly initialize their parameters.
The training follows~\citet{Liu:2020mbart} except that we tuned the learning rate within [3e-5,1e-3] and the dropout within [0.3,0.5] for the vanilla model and BT models.
We used the single model with the best validation perplexity for testing.
The length penalty is 1.0 and the beam size is 5.
We used sacreBLEU~\citep{post-2018-call} to calculate BLEU~\citep{papineni2002bleu} and TER~\citep{snover2006study} scores with the specific tokenization~\citep{Liu:2020mbart} for Romanian and the default tokenization for Russian.

\section{Understanding PT and BT}
\label{sec:understand}
In this section, we aim to better understand the similarities and differences between PT and BT on improving model performance.
We design two probing tasks to study the research question: {\em Which module of NMT do PT and BT respectively play a greater role in enhancing translation quality?}

\begin{table}[t]
\centering
\setlength{\tabcolsep}{5pt}
\begin{tabular}{cccrrr}
\toprule
\multirow{2}{*}{\bf Enc.}& \multirow{2}{*}{\bf Dec.}& \multicolumn{2}{c}{\bf PT} & \multicolumn{2}{c}{\bf BT} \\
\cmidrule(lr){3-4} \cmidrule(lr){5-6} 
 & &\bf BLEU &\bf $\Delta$  &\bf BLEU &\bf $\Delta$ \\
\midrule
N & N &33.7 &-    &33.7 &- \\ \hdashline 
N & Y &33.5 &-0.2    &\bf 37.8 &\bf +4.1 \\ 
Y & N &\bf 36.9 &\bf +3.2    &35.8 &+2.1 \\ \hdashline 
Y & Y &37.7 &+4.0    &38.3 &+4.6 \\   
\bottomrule
\end{tabular}
\caption{The probing tasks of PT and BT. NMT models are trained and evaluated on the WMT16 En-Ro benchmark. ``Y'' denotes the corresponding parameters are activated when augmented with PT or BT, while ``N'' denotes the inactive operation. PT and BT respectively contribute more to the NMT encoder and decoder.}
\label{tab:ablation}
\end{table}

\subsection{Effects of PT on NMT}
\label{sec:pteffect}
Given a pre-trained model, it is common to use its part or all parameters to initialize the downstream tasks. 
We design four NMT models, which differ from the NMT components (Encoder vs. Decoder) with parameter initialization manners (Random vs. Pre-trained).
As shown in Table~\ref{tab:ablation}, the variants are: 1) {\bf NN} is a vanilla NMT model, of which parameters are randomly initialized; 2) {\bf NY} means that parameters of NMT encoder are randomly initialized while those of decoder are initialized with pre-training; 3) {\bf YN} is contrary to {\bf NY}; and 4) {\bf YY} indicates that the whole NMT parameters are initialized with the pre-trained model. 
After that, the NMT models are fine-tuned on the parallel data with the same training strategy.

\paragraph{PT Mainly Contributes to Encoder}
As seen, the {\bf YY} initialization strategy significantly improves the vanilla NMT model by +4.0 BLEU scores, which reconfirms the effectiveness of PT for translation tasks~\cite{Liu:2020mbart}. By comparing {\bf NY} and {\bf YN}, we find that the pre-trained encoder can still help the NMT to achieve +3.2 BLEU improvements while the pre-trained decoder can only perform on par with the vanilla model (i.e. -0.2 BLEU). This demonstrates that PT mainly contributes to the encoder part of NMT model, and this claim is consistent with the conclusion with other pre-trained models. For instance, ~\citet{rothe2020leveraging} show that the NMT encoder initialization is superior to the decoder one when using pre-trained weights of BERT.
We hypothesize that the performance boost with PT mainly comes from the better ability of source-side understanding, which is significant to NMT such as on disambiguating word senses~\cite{tang2019encoders}.

\subsection{Effects of BT on NMT}
\label{sec:bteffect}
A vanilla NMT model is trained on the original bi-text and then fine-tuned on the mixture of the original and synthetic (i.e. back-translated) data.
We also design four NMT models, which differ from which parts of parameters are updated at the fine-tuning stage. 
As shown in Table~\ref{tab:ablation}, the variants are: 1) {\bf NN} is a vanilla NMT model only trained on the original data; 2) {\bf NY} indicates that parameters of the NMT encoder are fixed while those of decoder are updated during model fine-tuning; 3) {\bf YN} acts in an opposite way compared with {\bf NY}; 4) {\bf YY} means that the whole NMT parameters are updated at the fine-tuning stage. 

\paragraph{BT Mainly Contributes to Decoder}
BT has been sufficiently validated to improve the performance of NMT models ~\citep{edunov-etal-2018-understanding, edunov2020evaluation}. 
By exploiting additional target sentences, the NMT decoder can be enhanced to generate more fluent sentences in the target language. 
In contrast, the synthetic source sentences contain noises, which may be less useful for improving the ability of understanding. 
The results verify our hypothesis: BT mainly improves the decoder module of NMT. As seen, fine-tuning the whole NMT model (i.e. {\bf YY}) with BT data can gain the best performance (+4.6 BLEU than the vanilla model), which shows the effectiveness of BT method. 
Surprisingly, only fine-tuning the decoder (i.e. {\bf NY}) can perform close to {\bf YY} model (37.8 vs. 38.3 BLEU), which confirms our claims. Compared with {\bf NY}, the {\bf YN} model obtains relatively fewer improvements (+4.1 vs. +2.1 BLEU), showing that BT brings more benefits to the decoder than the encoder.

\section{Improving PT and BT}
The answer of the research question in Section~\ref{sec:understand} is: {\em PT and BT respectively contribute more to the NMT encoder and decoder, demonstrating that they are orthogonal and complementary to each other.}
This finding motivates us to better combine these two individual techniques together for further improving NMT models.

\subsection{Experiments}
As detailed in Section~\ref{sec:setup}, we conducted experiments on two commonly-used benchmarks En-Ro and En-Ru. Besides, we train the BT models from scratch instead of fine-tuning in Section~\ref{sec:bteffect}. As {\bf YY} models (in Table~\ref{tab:ablation}) always achieve best performances when augmented PT or BT, we update all parameters of NMT models in next experiments.

\begin{table}[t]
\centering
\setlength{\tabcolsep}{3pt}
\scalebox{0.92}{
\begin{tabular}{lrrrr}
\toprule
\multirow{2}{*}{\bf Model} & \multicolumn{2}{c}{\bf En-Ro} & \multicolumn{2}{c}{\bf En-Ru}  \\
\cmidrule(lr){2-3} \cmidrule(lr){4-5} 
& \bf BLEU &\bf TER &\bf BLEU &\bf TER \\
\midrule
\multicolumn{5}{c}{\it Existing Baselines} \\ 
XLM-R\scriptsize~\citep{conneau2019unsupervised} &35.6  &- &- &-\\
mRASP\scriptsize~\citep{lin2020pre} &37.6  &- &- &-\\
mBART\scriptsize~\citep{Liu:2020mbart} &37.7  &- &- &-\\ \midrule
\multicolumn{5}{c}{\it Our Implemented Systems} \\
Vanilla NMT &33.7 &48.6 & 28.8 &61.6  \\
~~+ PT &37.7 & 45.0 &31.6 &58.5 \\  \hdashline 
~~+ BT &38.4 & 45.0 &31.1 & 59.2 \\ 
~~+ BT + PT &41.2  & 42.6 &33.2  &57.1\\   \hdashline 
~~+ Tagged BT &38.6 & 44.9  &31.2 &59.3 \\ 
~~+ Tagged BT + PT &\bf 41.6 & \bf 42.1 &\bf 33.6 &\bf 56.5 \\   
\bottomrule
\end{tabular}}
\caption{Translation quality on the En-Ro and En-Ru benchmarks. ``+'' means incorporating PT and (Tagged) BT into NMT models. }
\label{tab:main}
\end{table}

The results are shown in Table~\ref{tab:main}. 
We use the vanilla model as our baselines, which are trained on original datasets with random initialization. 
Besides, we report results on existing PT models as our strong baselines, including XLM-R, mRASP, mBART.
As seen, PT can significantly improve the translation quality in all cases compared with vanilla baselines (averagely +2.5 BLEU), which is consistent with (or better than) existing PT models (37.7 vs. 35.6$\sim$37.7 BLEU). 
Furthermore, two BT methods\footnote{Tagged BT is to add a special token at the beginning of each back-translated source sentence.} (i.e. BT and Tagged BT) perform closely, which improves the standard NMT models by +3.5/+3.7 BLEU points on average. 
Simply combining them (+BT+PT) can further boost performances for NMT models across different sizes of datasets, showing the robustness and effectiveness of this approach.
Encouragingly, the combination of Tagged BT and PT performs better than the simple one, leading to state-of-the-art performances on the two benchmarks.
Similar tendencies are observed in terms of the TER scores.
The above results illustrate the better complementarity between PT and Tagged BT on improving translation quality for NMT models.

\begin{table}[t]
\centering
\begin{tabular}{lrrr}
\toprule
\bf Model & {\bf All} & {\bf ~~Src} & {\bf ~~Tgt} \\
\midrule
Vanilla &33.7 &29.4  &38.3 \\
~~+ PT &37.7 &33.8 &42.0 \\
\hdashline
~~+ BT &38.4 &31.5 &45.4 \\
~~+ BT + PT &41.2   &33.3 &48.6  \\ 
\hdashline 
~~+ Tagged BT & 38.6 &31.9 & 45.6 \\
~~+ Tagged BT + PT &\bf 41.6 &\bf 34.8  & \bf48.7 \\ 
\bottomrule
\end{tabular}
\caption{Translation quality of source-original and target-original sentences on the En-Ro benchmark. ``Src'' and ``Tgt'' respectively denote the sub-testsets of source-original and target-original while ``All'' means the whole testset.}
\label{tab:translationese}
\end{table}

\subsection{Analysis}
We conducted extensive analyses to better understand the improvement of our approach. All results are reported on the En-Ro benchmark.

\paragraph{Effects of Sentence Type}
Recent studies have shown that the evaluation of BT is sensitive to the sentences types, thus we report BLEU scores on the subsets of source-original (Src-Ori) and target-original (Tgt-Ori) datasets~\citep{zhang2019effect,liu-etal-2021-copying,wang-etal-2021-language-coverage}.\footnote{Src-Ori denotes the testing data originating in the source language, while Tgt-Ori denotes the data translating from the target language.}
Generally speaking, the translation of Src-Ori is more important than that of Tgt-Ori for practical NMT systems~\citep{graham-etal-2020-statistical}, thus its performance should be taken seriously.
As shown in Table~\ref{tab:translationese}, PT performs better on Src-Ori than BT (33.8 vs. 31.9 BLEU) while BT achieves higher scores on Tgt-Ori than PT (45.6 vs. 42.0 BLEU). 
Besides, simply combining PT and BT can improve the translation quality on both Src-Ori and Tgt-Ori sentences, but the improvement of Src-Ori is lower than only using PT.
By introducing tagged BT, the model can achieve better performance than the simple one, especially on source-original sentences. 
{\bf Takeaway:} {\em 1) PT and BT complementary in terms of originality of sentences; 2) Tagged BT can alleviate the bias of translating Tgt-Ori sentences which is significant to practical NMT systems.}

\begin{table}[t]
\centering
\begin{tabular}{lrrr}
\toprule
\bf Model & {\bf All} & {\bf Low} & {\bf High} \\
\midrule
Vanilla & 62.8 & 48.5 &64.6 \\
~~+ PT & 65.8 & 58.2 &66.7 \\
\hdashline
~~+ BT &65.9 &57.5 &67.1 \\
~~+ BT + PT &67.8 &60.8 & 68.8 \\ 
\hdashline 
~~+ Tagged BT & 66.1 &57.5 & 67.3 \\
~~+ Tagged BT + PT & \bf 68.3 &\bf 61.8 &\bf 69.1 \\ 
\bottomrule
\end{tabular}
\caption{F-measure of word translation according to frequency on the En-Ro benchmark. ``Low'' and ``High'' respectively denote the buckets of low- and high-frequency words while ``All'' means the whole words in the test set. Simply combining PT and BT improves the model performance, while adding tags to BT data further improves}
\label{tab:frequency}
\end{table}

\paragraph{Effects of Word Frequency}
Data augmentation is an effective way to improve the translation quality of low-frequency words~\citep{sennrich2016improving}. 
Thus, we compare the performance of the models on translating different frequencies of words. Specifically, we employed {\it compare-mt}~\citep{neubig-etal-2019-compare} to calculate the f-measure of translating low- and high-frequency words ($<$50 vs. $\geq$50). 
As shown in Table~\ref{tab:frequency}, PT improves more on translating low-frequency words (58.2 vs. 57.5 scores) while BT performs better on high-frequency words (67.3 vs. 66.7 scores). 
Furthermore, the combination of PT and tagged BT achieves the best performance on both low- and high-frequency words, leading to an overall improvement on the whole words.
Similar phenomenons can be observed by combining self-training and BT~\citep{ding-etal-2021-rejuvenating}. 
{\bf Takeaway:} {\em 1) PT and BT complementary in terms of frequency of words; 2) Tagged BT are more complementary to PT on lexical translation.}

\section{Conclusion and Future Works}
This paper broadens the understandings of pre-training (PT) and back-translation (BT).
We propose two probing tasks to investigate the impact of PT and BT on each NMT module and find that PT is more beneficial to the encoder while BT mainly improves the decoder.
Experimental results on the WMT16 English-Romanian and English-Russian benchmarks show that PT is nicely complementary to BT.
We also demonstrate that Tagged BT (i.e., adding tags to BT data) can further improve the complementarity of translating source-original sentences and low-frequency words.

In the future, we would like to apply curriculum learning~\citep{liu-etal-2020-norm,Zhan_Liu_Wong_Chao_2021,ding-etal-2021-progressive} to better organize the learning of PT and BT. It is also worthwhile to explore other kinds of methods utilizing monolingual data (e.g., self-training~\citep{zhang-zong-2016-exploiting,He2020Revisiting,jiao-etal-2021-self}) and validate the findings on practical NMT systems~\citep{transmart} and more generation tasks~\citep{liu2021understanding}.

\section*{Acknowledgement} 
This work was supported in part by the Science and Technology Development Fund, Macau SAR (Grant No. 0101/2019/A2), and the Multi-year Research Grant from the University of Macau (Grant No. MYRG2020-00054-FST). We thank the anonymous reviewers for their insightful comments.

\bibliography{emnlp2021}
\bibliographystyle{acl_natbib}
\end{document}